# Assessing the Limitations of Large Language Models in Clinical Fact Decomposition

### v1

@@ Abstract
Verifying factual claims is critical for using large language models (LLMs) in healthcare. Recent work has proposed fact decomposition, which uses LLMs to rewrite source text into concise sentences conveying a single piece of information, as an approach for fine-grained fact verification. Clinical documentation poses unique challenges for fact decomposition due to dense terminology and diverse note types. To explore these challenges, we present FactEHR, a dataset consisting of full document fact decompositions for 2,168 clinical notes spanning four types from three hospital systems. Our evaluation, including review by clinicians, highlights significant variability in the quality of fact decomposition for four commonly used LLMs, with some LLMs generating 2.6x more facts per sentence than others. The results underscore the need for better LLM capabilities to support factual verification in clinical text. To facilitate future research in this direction, we plan to release our code at https://github.com/som-shahlab/factehr.
Assessing the Limitations of Large Language Models in
Clinical Fact Decomposition
Monica Munnangi*1,2, Akshay Swaminathan*3, Jason Alan Fries*2, Jenelle Jindal4, Sanjana Narayanan2, Ivan Lopez3, Lucia Tu2, Philip Chung7, Jesutofunmi A. Omiye3,8, Mehr Kashyap3, Nigam Shah5,6,9 1Khoury College of Computer Sciences, Northeastern University; 2Center for Biomedical Informatics Research, Stanford University; 3Department of Biomedical Data Science∗∗; 4Stanford Health Care; 5Department of Medicine∗∗; 6Clinical Excellence Research Center∗∗; 7Department of Anesthesiology, Perioperative & Pain Medicine∗∗; 8Department of Dermatology∗∗; 9Technology and Digital Solutions, Stanford Health Care  munnangi.m@northeastern.edu, {akshaysw, jfries}@stanford.edu
**footnotetext: Equal contribution ****footnotetext: Stanford School of Medicine
@@ 1 Introduction
Figure 1: Fact decomposition in clinical text with LLMs requires handling diverse document types and complex medical jargon (colors highlight examples).
Verifying and attributing factual claims is essential for using large language models (LLMs) in healthcare. Evaluation strategies were proposed to measure summarization quality, ensure claims are properly grounded (attributed to specific text in the clinical note) (Bannur et al., 2024), text generated for documentation tasks must be grounded in data present in the patient’s EHR. , and reduce hallucinations and related errors (Zheng et al., 2023). A core component of these strategies is fact decomposition—an approach that paraphrases complex text into multiple, concise statements, each expressing a single piece of information (Fabbri et al., 2022; Chen et al., 2023; Min et al., 2023).
Fact decomposition can be used to verify facts against source documents using QA systems, textual entailment, or other model-based evaluation methods. Textual entailment, in particular, has proven effective for automating factuality assessment (Kamoi et al., 2023; Ru et al., 2024; Xie et al., 2024). Although these verification techniques are well-studied in both general and scientific texts (Wadden et al., 2020; Wright et al., 2022), their application to clinical text remains underexplored. This represents a significant gap, as performance metrics are highly sensitive to the quality of fact decomposition (Wanner et al., 2024).
@@
The decomposition and verification of the facts in clinical notes using LLMs presents significant challenges. Clinical notes employ medical terminology, special acronyms, and non-grammatical shorthand (see Figure 1). Clinical observations are often compositional, as in: “lobulation at the apex of the left hemithorax along the mediastinal border is residual of slowly resolving hematoma." For research purposes, clinical notes are frequently stripped of their original markup—like tables, lists, and section headers—thereby increasing ambiguity and parsing difficulties. The de-identification process, which masks patient names and dates, further introduces noise that can confound LLMs. Moreover, medical documentation serves diverse functions: nursing notes offer brief, time-stamped updates on a patient’s physiological status, whereas discharge summaries integrate an entire hospital stay’s details for billing and compliance purposes. Consequently, these documents differ substantially in presentation, length, fact density, and temporal scope.
@@
Existing entailment datasets – which denote whether a piece of text (the hypothesis) logically follows from another piece of text (the premise) – do not capture the diversity of clinical documentation, either because they focus on a single note type (Miura et al., 2021), use artificial data (Romanov and Shivade, 2018a), or have a limited representation of hospital systems. Fact decomposition across different types of clinical documents has not been evaluated. In addition, it is frequently unclear how to evaluate the quality of a decomposition because determining entailment requires medical expertise. Finally, frontier model APIs are often not HIPAA-compliant, hindering evaluations using popular datasets such as MIMIC (Johnson et al., 2016). Given these challenges, we explore the following research questions (RQ’s):
RQ1 — Fact decomposition
@@
How do the length, quality, and similarity of fact decompositions differ across LLMs?
RQ2 — Entailment
How accurately can LLMs determine textual entailment of clinical documents?
RQ3 — Fact verification
How accurate and complete are the fact decompositions of clinical notes created by LLMs?
We present FactEHR, a fact decomposition and textual entailment dataset derived from 2,168 clinical notes across four document types from three hospital systems. Using FactEHR, we contribute the following:
  1. 1.
    For RQ1, we analyze the length and similarity of fact decompositions from four LLMs and conduct a qualitative evaluation with clinical experts.
  2. 2.
    For RQ2, we compare LLMs in their ability to perform textual entailment using standard NLI datasets and expert-labeled entailment pairs from clinical documents. We establish feasibility of using GPT-4o to assess entailment.
  3. 3.
    For RQ3, we quantify the accuracy and completeness of fact decompositions from four LLMs.
Our findings reveal considerable variation in the number and atomicity of fact decompositions across clinical documents and across LLMs, with some LLMs generating 2.6x more facts per sentence than others. This raises questions about the validity of metrics that rely on fact decompositions for evaluating LLMs in healthcare documentation tasks.
@@ 2.1 Experimental Setup
We use LLMs to perform fact decomposition—the process of breaking down source text into concise, atomic statements that convey individual pieces of information. We then examine the resulting fact decompositions, comparing their lengths and similarity across different LLMs. Finally, we pair each decomposition with its corresponding source note to form entailment pairs, evaluating how accurately and completely LLMs capture the original information.
@@
We consider four LLMs to perform fact decomposition: GPT-4o (OpenAI, 2024), o1-mini (Singhal et al., 2023b), Gemini-1.5-Flash-002 (et al., 2024), and Llama3-8b-Instruct (Dubey et al., 2024). We will refer to Llama3-8b-Instruct as Llama3 or Llama3-8B and Gemini-1.5-Flash-002 as Gemini-1.5 in the rest of the paper (unless specified otherwise). All models were run using HIPAA-compliant compute environments and APIs. Model generation hyperparameters are included in Appendix C.
@@
Adapted from (Min et al., 2023), our prompt includes two in-context examples of clinical note fact decompositions and instructs the LLM to output independent facts as a delimited string (Appendix 11).
@@ 2.3 Note Types
We consider four clinical note types: (1) Procedure Note, which typically includes the procedures, clinical indications, findings, and follow-up recommendations. (2) Nursing Note, which provides a systematic assessment of a patient’s condition across body systems (e.g., cardiovascular, neurologic) at a specific time, with less emphasis on future care planning. (3) Progress Note, which summarizes a patient’s medical status from the previous day and outline the care plan for the next. (4) Discharge Summary, which provides a concise overview of the patient’s presentation, past medical history, key findings, future medical plans, and discharge medications for subsequent care providers.
2.4 RQ1 — Fact Decomposition
We examine the length and similarity of fact decompositions across LLMs, as well as the atomicity of source sentences. Figure 2 outlines example decompositions across all LLMs used in this study.
Fact Counts
Fact Similarities
Sentence Atomicity
2.5 RQ2 — Entailment
To evaluate the accuracy and completeness of fact decompositions, we pair each decomposition with its corresponding source note to form entailment pairs. We investigate whether LLMs can accurately evaluate textual entailment in this setting.
@@ 2.5 RQ2 — Entailment
We first benchmark five LLMs—GPT-4o, GPT-4o-mini, Gemini-1.5, Llama3-8B, and Llama3-70B —using four publicly available natural language inference (NLI) benchmarks—SciTail (Khot et al., 2018), MedNLI (Romanov and Shivade, 2018a), MultiNLI (Williams et al., 2018a), and SNLI (Bowman et al., 2015) (Appendix A.2)—and the FactEHR development set (described below).
We tune our prompt using 40 premise-hypothesis pairs sampled from these datasets, optimizing for F1 score. The final prompt, adapted from (Xie et al., 2024), instructs the LLM to output a binary entailment judgement in JSON format (Appendix 12).
We then selected the model with the highest F1 score and applied it to all 987,266 entailment pairs in the FactEHR dataset (details provided below). Of these pairs, 1,036 were manually annotated by clinical experts, allowing us to directly assess the LLM’s accuracy on the FactEHR dataset.
2.6 RQ3 — Fact Verification
@@
To assess the accuracy and completeness of model-generated fact decompositions, we use fact-precision and fact-recall, entailment-based metrics proposed by (Xie et al., 2024).
@@
Let $S$ denote the set of sentences obtained by tokenizing the clinical note $d$, and let $C$ represent the set of factual claims produced by decomposing $d$. We define $[d\models c]=1$ if a hypothesis $c$ is completely entailed by a premise $d$, and 00 otherwise. For example, the hypothesis, “There is a pleural effusion on the left side,” is completely entailed by the premise “Left pleural effusion increased from prior scan.”
@@
where $S$ is the set of sentences in the clinical note, and $[C\models s]$ is an indicator function that evaluates to 1 if $s$ is entailed by $C$, and 0 otherwise.
@@
3 Results
@@
From these 987,266, we randomly selected 1,036 pairs for manual entailment annotation by clinical experts. Clinical experts included two board certified physicians, two residents, two medical students, and a clinical researcher. Annotators were provided instructions for labeling entailment (Appendix D), and a subset of 100 entailment pairs were labeled in duplicate to calculate inter-rater agreement.
Note Type Model Name Mean $\pm$ SE
Procedure Note GPT-4o 1.45 $\pm$ 0.03
o1-mini 1.45 $\pm$ 0.03
Gemini-1.5 1.12 $\pm$ 0.03
Llama3 8B 0.89 $\pm$ 0.03
Nursing Note GPT-4o 2.37 $\pm$ 0.14
o1-mini 2.42 $\pm$ 0.13
Gemini-1.5 2.05 $\pm$ 0.12
Llama3 8B 1.58 $\pm$ 0.12
Progress Note GPT-4o 1.28 $\pm$ 0.03
o1-mini 1.46 $\pm$ 0.03
Gemini-1.5 1.02 $\pm$ 0.02
Llama3 8B 0.73 $\pm$ 0.04
Discharge Summary GPT-4o 1.30 $\pm$ 0.02
o1-mini 1.55 $\pm$ 0.03
Gemini-1.5 0.98 $\pm$ 0.02
Llama3 8B 0.60 $\pm$ 0.04
Table 1: Average facts per sentence by note type. Red indicates less facts than sentences meaning the LLM is likely failing to capture all facts.

### v2

@@ Abstract
Verifying and attributing factual claims is essential for the safe and effective use of large language models (LLMs) in healthcare. A core component of factuality evaluation is fact decomposition, the process of breaking down complex clinical statements into fine-grained atomic facts for verification. Recent work has proposed fact decomposition, which uses LLMs to rewrite source text into concise sentences conveying a single piece of information, to facilitate fine-grained fact verification. However, clinical documentation poses unique challenges for fact decomposition due to dense terminology and diverse note types and remains understudied. To address this gap and explore these challenges, we present FactEHR, an NLI dataset consisting of document fact decompositions for 2,168 clinical notes spanning four types from three hospital systems, resulting in 987,266 entailment pairs. We assess the generated facts on different axes, from entailment evaluation of LLMs to a qualitative analysis. Our evaluation, including review by the clinicians, reveals substantial variability in LLM performance for fact decomposition. For example, Gemini-1.5-Flash consistently generates relevant and accurate facts, while Llama-3 8B produces fewer and less consistent outputs. The results underscore the need for better LLM capabilities to support factual verification in clinical text.
@@ 1 Introduction
Verifying and attributing factual claims is essential for the safe and effective use of large language models (LLMs) in healthcare. Evaluation strategies have been proposed to assess summarization quality, ensure claims are properly grounded (i.e., attributed to specific text in the clinical note), and reduce hallucinations and related errors (Zheng et al., 2023). A core component of these strategies is fact decomposition, which rewrites complex text into concise, atomic statements, each conveying a single piece of information(Fabbri et al., 2022; Chen et al., 2023; Min et al., 2023).
Fact decomposition is used to verify facts against source documents using QA systems, textual entailment, or other model-based evaluation methods. Textual entailment, in particular, has proven effective for automating factuality assessment (Kamoi et al., 2023; Ru et al., 2024; Xie et al., 2024). Although these verification techniques are well studied in both general and scientific texts (Wadden et al., 2020; Wright et al., 2022), their application to clinical texts (Xie et al., 2024) remains underexplored. This represents a significant gap, as performance metrics are highly sensitive to the quality of fact decomposition (Wanner et al., 2024).
@@
The decomposition and verification of the facts in clinical notes using LLMs presents significant challenges. Clinical notes employ medical terminology, special acronyms, and non-grammatical shorthand. Clinical observations are often compositional, as in: “lobulation at the apex of the left hemithorax along the mediastinal border is residual of slowly resolving hematoma.” For research purposes, clinical notes are frequently stripped of their original markup—like tables, lists, and section headers—thereby increasing ambiguity and parsing difficulties. The de-identification process, which masks patient names and dates, further introduces noise that can confound LLMs. Moreover, medical documentation serves diverse functions: nursing notes offer brief, time-stamped updates on a patient’s physiological status, whereas discharge summaries integrate an entire hospital stay’s details for billing and compliance purposes. Consequently, these documents differ substantially in presentation, length, fact density, and temporal scope.
Figure 1: End-to-end pipeline of our work. We generate fact decompositions of clinical notes with LLMs. Evaluate the generated dataset in an entailment setting; For calculating Fact-Precision, we check if the generated fact (the hypothesis) logically follows from original clinical note (the premise) and vice versa for Fact-Recall. We validate the LLM-as-a-Judge with a subset of human evaluation. We also perform a modest amount of qualitative evaluation with experts.
@@
Existing entailment datasets – which denote whether a piece of text (the hypothesis) logically follows from another piece of text (the premise) – do not capture the diversity of clinical documentation, either because they focus on a single note type (Miura et al., 2021), use artificial data (Romanov and Shivade, 2018a), or have a limited representation of hospital systems. Fact decomposition across different types of clinical documents has not been evaluated. In addition, it is frequently unclear how to evaluate the quality of a decomposition because determining entailment requires medical expertise. Finally, frontier model APIs are often not HIPAA-compliant, hindering evaluations using popular datasets such as MIMIC (Johnson et al., 2016).
To address this gap and to explore these challenges, we present FactEHR, a fact decomposition and textual entailment dataset derived from 2,168 clinical notes across four document types from three hospital systems. In Figure 1 provides an overview of the pipeline used to generate fact decompositions, evaluate them using LLM-as-a-Judge (a method that uses a large language model to assess whether a claim is entailed, contradicted, or unsupported by a source text), and validate the results against a subset of expert evaluations. We also include a small set of qualitative human evaluations. Using this dataset, we explore the following questions and contribute toward answering them:
RQ1: Fact Decomposition
@@
How do fact decompositions vary in length, quality, and similarity across LLMs? We compare outputs from four models and analyze key differences.
RQ2: Entailment
How well can LLMs assess textual entailment in clinical documents? We evaluate performance on existing NLI benchmarks and a new set of expert-labeled entailment pairs derived from clinical notes.
RQ3: Fact Verification
How accurate and complete are LLM-generated fact decompositions of clinical notes? We assess performance using formal metrics and detailed clinical expert review.
Our findings reveal considerable variation in the number and atomicity of fact decompositions across clinical documents and across LLMs, with some LLMs generating 2.6x more facts per sentence than others. The quality of fact decompositions also vary highly between LLMs, especially closed models generating factually consistent and covering most information as opposed to smaller open source models, where they are incomplete or they get into infinite generation loops. Consequently, higher fact counts do not correspond to better coverage or consistency—smaller open models often omitted key clinical details or hallucinated, while closed models more reliably captured core information. This raises questions about the validity of metrics that rely on fact decompositions for evaluating LLMs in healthcare documentation tasks. To facilitate future research in this direction, we release code [^1] https://github.com/som-shahlab/factehr, and data [^2] https://som-shahlab.github.io/factehr-website/.
Generalizable Insights about Machine Learning in the Context of Healthcare
Our study highlights that fact decomposition is a crucial yet overlooked factor in evaluating LLMs for healthcare applications, as variations in decomposition quality significantly impact factuality assessments. We demonstrate that textual entailment, a key method for fact verification, faces unique challenges in clinical settings due to domain-specific complexities like shorthand, compositional observations, and structural inconsistencies. Additionally, we find substantial variability in how different LLMs generate fact decompositions, affecting downstream evaluations and raising concerns about the reliability of existing benchmarks. These insights underscore the need for standardized fact decomposition methodologies and more robust, domain-specific evaluation strategies to ensure accurate and trustworthy use of LLMs in clinical NLP tasks.
We release a natural language inference (NLI) dataset generated from real clinical notes spanning multiple note types and sources. Fact decompositions are produced by LLMs, with a subset evaluated by medical experts and scaled evaluation provided by LLM-as-a-Judge. This dataset can be used as (a) a benchmark for evaluating models on the fact decomposition task and (b) training models using both LLM-generated decompositions and their corresponding evaluations.
2 Related Work
Large Language Models (LLMs) in Healthcare
LLMs encode extensive clinical and biomedical knowledge, recent models specifically, GPT-4 (OpenAI, 2023), as well as Gemini (Team et al., 2024), have emerged as state-of-the-art LLMs, showcasing impressive capabilities in a wide range of domain-specific applications (Singhal et al., 2023a; Agrawal et al., 2022; Munnangi et al., 2024). Several evaluations have shown that LLMs can achieve performance comparable to fully supervised models on tasks such as entity extraction and relation extraction (Wadhwa et al., 2023), under both few-shot and zero-shot settings. While considerable progress has been made on atomic fact generation in general domains (Min et al., 2023; Gunjal and Durrett, 2024) and biomedical literature (Wadden et al., 2020; Wright et al., 2022), there is little research focusing on clinical notes. Existing frameworks such as MAIRA 2 (Bannur et al., 2024) address fact-checking in radiology reports but not the broader range of clinical documents.
Clinical NLI Datasets
Several related NLI datasets have been proposed in the clinical domain. The NLI4CT dataset focuses on clinical trial reports (Jullien et al., 2023) and represents one of the earliest efforts to apply NLI to scientific text, though it is limited in size, containing only 2,400 instances. Another dataset targets social determinants of health extracted from clinical notes (Lelkes et al., 2023). Our work differs substantially from both in terms of scale and the diversity of EHR note sources.
LLM-as-a-Judge
Evaluating language model outputs remains an open research challenge, particularly for generative tasks where multiple valid responses exist (Chang et al., 2023). In closed-world tasks with a small, discrete set of correct answers—such as classification or multiple-choice question answering—models typically use exact match or likelihood-based selection to produce outputs, which are then evaluated using standard metrics like accuracy. For short-form text generation, token-overlap metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) are commonly used. However, these metrics often fail to reward correct but paraphrased outputs. Semantic similarity metrics like BERTScore (Zhang et al., 2020) address this limitation by comparing contextualized token embeddings to measure some aspects of meaning, rather than relying on exact lexical overlap.
For complex generative tasks, evaluation is challenging due to the large space of plausible outputs and the potential mixing of correct and incorrect information within a single response. Recent work (Min et al., 2023; Xie et al., 2024; Tian et al., 2023) addresses this by decomposing text into atomic facts to assess factuality and completeness; for example, (Xie et al., 2024) compute claim-level precision and recall using entailment models. However, current models often struggle to generate atomic facts that are both accurate and comprehensive, especially in the clinical domain. In our experiments, entailment-based evaluation using LLMs showed strong correlation with human judgments. Given the high cost and limited scalability of expert annotation, we adopt an LLM-as-a-Judge framework—validated against human evaluations—to assess our dataset.
@@ 3.1 Experimental Setup
We use LLMs to perform fact decomposition, the process of breaking down source text into concise, atomic statements that convey individual pieces of information. We then examine the resulting fact decompositions, comparing their characteristics such as number of decomposed facts, length of the generated facts and similarity across different LLMs. Finally, we pair each decomposition with its corresponding source note to form entailment pairs, evaluating how accurately and completely LLMs capture the original information. In this section, we describe the models, prompts, datasets, types of clinical notes and evaluation, in depth, with more information in the appendix as appropriate.
@@
We consider four LLMs to perform fact decomposition: GPT-4o (OpenAI, 2024), o1-mini (Singhal et al., 2023b), Gemini-1.5-Flash-002 (Team et al., 2024), and Llama3-8b-Instruct (Dubey et al., 2024). For brevity, we will refer to Llama3-8b-Instruct as Llama3 or Llama3-8B and Gemini-1.5-Flash-002 as Gemini-1.5 in the rest of the paper (unless specified otherwise). We also conducted preliminary experiments with domain specific models including Google’s MedLM but found its performance subpar [^3] More information in Appendix C., and so did not pursue further. All models were run using HIPAA-compliant compute environments and APIs. Model generation hyperparameters are included in Appendix D.
@@
We use two distinct prompts: one for fact decomposition and another for entailment evaluation. The fact decomposition prompt, adapted from (Min et al., 2023), includes two in-context examples of clinical note fact decompositions and instructs the LLM to output independent facts as a delimited string (Appendix L, Figure 11). The entailment evaluation prompt is tuned using 40 premise-hypothesis pairs sampled from our datasets, optimizing for F1 score. The final version, adapted from (Xie et al., 2024), instructs the LLM to produce a binary entailment judgment in JSON format (Appendix 12).
@@ 3.3 Note Types
We consider four clinical note types: (1) Procedure Note, which typically includes procedures, clinical indications, findings, and follow-up recommendations. (2) Nursing Note, which provides a systematic assessment of a patient’s condition across body systems (e.g., cardiovascular, neurologic) at a specific time, with less emphasis on future care planning. (3) Progress Note, which summarizes a patient’s medical status from the previous day and outline the care plan for the next. (4) Discharge Summary, which provides a concise overview of the patient’s presentation, past medical history, key findings, future medical plans, and discharge medications for subsequent care providers. We provide more details about note types in Appendix B.
3.4 Entailment evaluation (Validating LLM-as-a-Judge)
This study benchmarks the LLM-as-a-judge approach for fact verification in clinical texts, a previously under-evaluated component. To ensure statistical validity of performance metrics like sensitivity and specificity, power calculations (Buderer, 1996) guided the scale of human annotations based on NLI results from the FactEHR dataset. We annotated 1000 unique entailment pairs (250 per note type) and included 200 duplicates (50 per note type) to assess inter-annotator agreement, totaling 1200 annotations in this tranche—sufficient for high-confidence estimates (99% overall,  80% per note type). Combined with the 2468 previously annotated pairs, our dataset now comprises  3500 human-annotated clinical entailment pairs, making it one of the largest of its kind in the domain.
To evaluate the accuracy and completeness of fact decompositions, we pair each decomposition with its corresponding source note to form premise–hypothesis pairs for textual entailment. We assess whether LLMs can reliably evaluate entailment in this setting (see Appendix I, Table 9 for details). As part of early model development, we also benchmarked prior models such as RoBERTa, finding that their performance was consistently lower than that of frontier LLMs like GPT-4o. Additionally, we explored entity-level metrics such as RadGraph F1 (Jain et al., 2021), but found they did not correlate well with expert judgments of factuality in our domain, highlighting the need for more nuanced and holistic evaluation approaches.
@@ 3.4 Entailment evaluation (Validating LLM-as-a-Judge)
We first benchmark five LLMs (GPT-4o, GPT-4o-mini, Gemini-1.5, Llama3-8B, and Llama3-70B) using four publicly available NLI benchmarks: SciTail (Khot et al., 2018), MedNLI (Romanov and Shivade, 2018a), MultiNLI (Williams et al., 2018), and SNLI (Bowman et al., 2015). Additional information about these datasets is provided in Appendix A.2. We also evaluate performance on the FactEHR development set, which is described below. All models use the entailment prompt shown in Figure 12, applied to all 987,266 entailment pairs in the FactEHR dataset. Of these, 1,036 pairs were manually annotated by clinical experts, enabling direct assessment of model accuracy.
3.5 Fact Verification
@@
To assess the accuracy and completeness of model-generated fact decompositions, we use fact-precision and fact-recall, entailment-based metrics proposed by DocLens (Xie et al., 2024).
@@
Let $S$ denote the set of sentences obtained by tokenizing the clinical note $d$, and let $C$ represent the set of factual claims produced by decomposing $d$. We define $d\models c]=1$ if a hypothesis $c$ is completely entailed by a premise $d$, and 0 otherwise. For example, the hypothesis, “There is a pleural effusion on the left side,” is completely entailed by the premise “Left pleural effusion increased from prior scan.”
@@
where $S$ is the set of sentences in the clinical note, and $C\models s]$ is an indicator function that evaluates to 1 if $s$ is entailed by $C$, and 0 otherwise.
@@
4 FactEHR
Complete details on data sources and preprocessing are provided in Subsection 3.2. Figure 2 shows example fact decompositions from each LLM evaluated in this study. Table 1 enumerates key dataset summary statistics.
4.1 Data Generation
@@
Name Creation Rule Size Description
FactNotes note $\in$ D 2,168 A collection of clinical notes from the dataset D.
FactDecomp note $\rightarrow$ fact-list 8,665 Decomposition of notes into a list of structured, atomic facts.
FactEntail_p I[note $\Rightarrow$ fact] 491,663 Entailment pairs for evaluating whether a note implies a given fact.
FactEntail_r I[fact-list $\Rightarrow$ sentence] 495,603 Entailment pairs for evaluating whether a fact list implies a full sentence.
FactEntail FactEntail_p $\cup$ FactEntail_r 987,266 Union of both entailment settings to support broad NLI evaluation.
FactEntail_ann x $\sim$ U(FactEntail) 1,036 Human-annotated entailment data sampled from FactEntail_p and FactEntail_r.
Table 1: FactEHR dataset overview. We generate fact decompositions from 2,618 clinical notes and derive multiple NLI subsets.
@@
From these 987,266, we randomly selected 1,036 pairs for manual entailment annotation by clinical experts. Clinical experts included two board certified physicians, two residents, two medical students, and a clinical researcher. Annotators were provided instructions for labeling entailment (Appendix E), and a subset of 100 entailment pairs were labeled in duplicate to calculate inter-rater agreement.
5 Results